OpenStudy (anonymous):

I'm trying to solve some differential equations, and I think I have the right answer, but I'm not sure :( With two equations.

OpenStudy (anonymous):

These are the two equations I'm dealing with. I think I have the first one so far.

OpenStudy (anonymous):

what type quadratic, simulatneous?

OpenStudy (anonymous):

\[\frac{ dC }{ dt }=-k _{1}C \]

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ dC }{ C } = \frac{ -k _{1} Cdt}{ C}\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ dC }{ C }=\int\limits_{}^{}-k _{1}dt\]

OpenStudy (anonymous):

\[\ln (C)=-k _{1}\int\limits_{}^{}dt\]

OpenStudy (anonymous):

\[\ln (C)=-k _{1}t+C _{1}(constant of integration)\]

OpenStudy (anonymous):

I'm not sure how to continue from here, or if this is right :|

OpenStudy (amistre64):

e^(lnx) = x

OpenStudy (amistre64):

\[\ln (C)=-k _{1}t+\delta\] \[exp(\ln (C))=exp(-k _{1}t+\delta)\] \[C=exp(-k _{1}t+\delta)\] \[C=exp(-k _{1}t)~exp(\delta)\] \[C=\lambda~exp(-k _{1}t)\]

OpenStudy (amistre64):

when t=0, C = lambda = Ao to satisfy your initial condition therefore: \[\large C=A_oe^{-k_1t}\]

OpenStudy (amistre64):

\[\large \frac{dG}{dt}=k_1~A_oe^{-k_1t}-k_2G+k_2G_o\] that last term is a constant \[\large G'+k_2G=k_1~A_oe^{-k_1t}+k_2G_o\]

OpenStudy (amistre64):

any of this make sense?

OpenStudy (anonymous):

Hey! Thank you so much for the help, I'm sorry I didn't get back to you sooner, I was in class. But yes, that makes sense to me! We didn't really go over much of this in class so I was having trouble getting it. Thanks again!!